OpenStudy (anonymous):

Linear Equation in 3 Unknown. Using 2 Methods. A. By Elimination B. Cramer's Rule.

OpenStudy (anonymous):

Equations: 3x-2y+3z=2 4x+3y-2z=3 2x-4y+3z=5

OpenStudy (anonymous):

Normally the managment of this site does not approve of answers only being posted, however, this problem requires two solving methods and so the solution process for each method is expected to be shown. The thing is is that the solutions for x, y and z are not integer numbers; they are fractions as shown below:\[\left\{x=\frac{41}{31},y=-\frac{67}{31},z=-\frac{65}{31}\right\} \]